I need to know step by step on what percent of $84 is of $10.50. I have the answer which is 12.5 but what are the steps in getting that answer.
Answer by stanbon(75887)

I need to know step by step on what percent of $84 is of $10.50. I have the answer which is 12.5 but what are the steps in getting that answer.
-----
Here's one way to do it.
Percent means per hundred.
x/100 = 10.50/84
---
x = 100(10.50/84)
---
x = 1050/84
--
x = 12.5
---
Ans: 12.5%
==================
Cheers,
Stan H.